# Hamiltonian replica exchange augmented with diffusion-based generative models and importance sampling to assess biomolecular conformational basins and barriers

### v1

@@ 2.7 Comparison with TREM and US
We now provide a critical comparison between the results of DDPM-REST2 and that of DDPM-TREM using a similar simulation setup. We performed TREM simulations using 32 replicas of 600 ns ranging from 279 K to 600 K, yielding an exchange rate of 0.13 (this illustrates the much lower performance of T-REM compared to REST2 despite three times more replicas).
Figure 3: Comparison of two-dimensional free-energy surfaces computed along RMSD and $R_{g}$, obtained from TREM (A), REST2 (B), TREM refined by DDPM (C) and REST2 refined by DDPM (D), as well as the free-energy values along the minimum free-energy path as predicted by DDPM (E). White lines on (C) and (D) show the minimum free-energy path. (F) One-dimensional free energy profile computed along the RMSD direction, comparing the training data from TREM and REST2 (dashed lines, orange and blue, respectively) with the DDPM predictions (solid lines, orange and blue, respectively), and Umbrella Sampling as a benchmark (solid green line). Representative structures of basins $\alpha$, $\beta$, $\gamma$, and $\delta$ are shown in Figure S1.
By projecting the probability distributions estimated from the unscaled replica along (Figure 3A and B), it is clear that both T-REM and REST2 achieve comparable sampling performances. Key free-energy minima are identified, corresponding to radically different conformations (a few examples numbered A to D are shown in Figure S1). We observe that while the basins corresponding to regions A and B are well sampled, there are significantly fewer samples in C and D, and the transition regions are not well refined. As a consequence, once cannot clearly estimate the minimum free-energy paths and barriers connecting the apparent basins, in particular $\alpha$, $\beta$ and $\gamma$ on this two-dimensional surface.
When post-processing the results of the simulations with DDPM, and then use it to generate new samples for the unscaled replica conditions, much more converged free-energy surfaces can be recovered (Figure 3C and D). In particular, the transition states regions connection $\alpha$, $\beta$ and $\gamma$ are now fully resolved. These are extrapolated from the model based on higher temperature/lower potential simulations learned from the model and extrapolated in the unperturbed replica. Indeed, the corresponding areas appear sampled when climbing the replica ladder (Figure S2).
@@
TREM TREM+DDPM REST2 REST2+DDPM
$F_{\beta}-F_{\alpha}(kcal/mol)$ 1.11 $\pm$ 0.34 1.050 $\pm$ 0.002 1.23 $\pm$ 0.24 1.100 $\pm$ 0.002
$F_{\gamma}-F_{\beta}(kcal/mol)$ 0.81 $\pm$ 0.36 0.720 $\pm$ 0.002 0.64 $\pm$ 0.30 0.81 $\pm$ 0.05
Table 1: Total free energy differences between basins A, B, and C computed from TREM, TREM+DDPM, REST2 and REST2+DDPM. Error bars for the TREM and REST2 simulations were obtained by calculating standard deviations across blocks of 100 ns. For DDPM, error bars correspond to the standard deviation calculated across four sets of 1000000 generated points, obtained from the model trained on the same dataset.
Although there is good agreement between the two replica exchange strategies combined with DDPM, there is no guarantee that they actually converge toward the actual ground-truth reference of the associated conformational change. Having determined from 2D-projections that the key bottleneck to conformational exploration was (in the chosen CV-space) transitions along the RMSD to the crystal structure, we also performed US simulations along this coordinate to estimate a ground-truth 1D FES to which we could compare the results of replica exchange simulations augmented with DDPM.
For RMSD values below 3 Å, there is good agreement between the DDPM predictions trained on both TREM and REST2, and the umbrella sampling simulations. For RMSD larger than 3 Å, the agreement is less accurate. This is due to the fact that high RMSD values correspond to a large number of possible unfolded, disordered conformations of CLN025 that are difficult to sample in all approaches (US simulations are also unlikely to be converged in that region). In particular, this indicates that the sampling provided by the replicas is not fully converged in this region of phase space. In particular, sampling is pretty sparse and noisy in these high-RMSD regions, and even so for the highest temperature replica (Figure S2).
@@ 2.8 Importance of exchanges for reliable reference data
The fact that DDPM results remain sensitive to the quality of sampling in the replica exchange simulations is further illustrated by studying the impact of exchanges among replicas on the obtained results. Strikingly, it has been argued in the original DDPM-TREM formulation22, as well as in subsequent work26, that DDPM could perform well even if exchanges were vanishingly small (smaller than a few %, or even non existent).
@@
As seen in these plots, exchanges are crucial for converging the melting curve, which is a measure of the convergence of the thermodynamic quantities, and they significantly impact DDPM predictions. For instance, DDPM trained on TREM without exchanges, starting with unfolded structures, cannot generate folded structures because no folding is observed in the TREM without exchanges. Starting from a folded structure, TREM simulations can generate unfolded structures in the high-temperature replicas, but the melting curves are very noisy and DDPM does not improve the sampling alone.
@@
The main difference with previous work is that (i) here exchanges are crucial to drive chignolin folding/unfolding, whereas some of the systems investigated previously did not exhibit folded/unfolded states22; and (ii) in other more complex cases26, configurations are mixed across the replica ladder and periodically updated, therefore mitigating the effects of small exchange rates.
This highlights the advantage of using REST2 for training DDPM, as REST2 requires much fewer replicas compared to T-REM. Regarding computational cost, the training of DDPM and data generation for REST2 and TREM in these examples were similarly affordable from a computational perspective (less than a dozen gpu.hrs). The gain in computational efficiency comes from the better scaling of REST2 with system size: we needed only 12 replicas with REST2, compared to 32 with TREM, achieving better exchange rates with REST2 (0.51) than with TREM (0.13) and yielding comparable sampling at the reference temperature (Figure 3A). This resulted in a 2.7-fold reduction in computational cost for REST2 compared to TREM for the same sampling efficiency. The gain would be even more obvious for a larger biomolecule that would require many more replicas in TREM.
@@ 2.9 Resolving large conformational changes with high activation barriers
We thus chose to focus on the loop opening/closing motion in the phosphatase PTP1B. PTP1B is a well-studied member of the Protein Tyrosine Phosphatase (PTP) superfamily, which is involved in multiple cellular processes such as glucose uptake, and proliferation 45. It is also a drug target for several diseases, including diabetes and cancer 46.
@@
Before we focus on our results, we now provide some context about this enzyme (Figure 5A). PTP1B catalyzes the hydrolysis of phosphorylated tyrosine in a two-step process involving a thiophosphate enzyme intermediate. The first step consists of a nucleophilic attack by a cysteine residue on the phosphorylated tyrosine, breaking the phosphate bond. The second step involves the hydrolysis of the intermediate through a nucleophilic attack by a water molecule. Both steps require the assistance of an aspartic residue (Asp181) which acts as a proton donor and acceptor during the reaction. Asp181 belongs to the WPD-loop, a flexible loop comprising a dozen residues (177 to 188) that exists in two conformations: the hydrolysis-incompetent open state, and the hydrolysis-competent closed state (Figure 5A). In the open state, Asp181 forms a salt-bridge with arginine Arg112, while in the closed state, Asp181 rotates to form a salt-bridge with Arg221, allowing Asp181 to get closer to the catalytic site for proton transfers.
@@
Starting first from a closed conformation (we provide further below the comparison with simulations starting from the open conformation), we propagated the simulations for 600 ns and analyzed the converged part of the trajectories. We computed the free-energy surface along the two torsional angles $\psi_{181}$ and $\phi_{182}$ whose changes allow the rotation of the Asp181-Phe182 peptide group, which have been shown to be responsible for the high free-energy barrier characterizing the transition 31. The FES (Figure 6A) is characterized by two basins, corresponding to the closed (upper left) and open (bottom right) conformations; no transition path is sampled to go from the closed to the open conformation. In addition, the free-energy difference between the two basins is in favor of the closed conformation, at odds with previous investigations and experimental results (Table 2).
(kcal/mol) R2 (c) R2-DDPM (c) R2-b (c) R2-b-DDPM (c) R2-b-DDPM (o)
$\Delta F_{\text{o-c}}$ 0.70 $\pm$ 0.24 0.870 $\pm$ 0.001 -0.80 $\pm$ 1.08 -0.920 $\pm$ 0.003 -1.750 $\pm$ 0.002
Table 2: Free energy differences between closed (c) and open (o) conformations, computed from REST2 (R2 in the table) starting closed, REST2+DDPM, REST2+Bias starting closed, REST2+Bias+DDPM, and REST2+DDPM+Bias starting open. The experimental value for $\Delta F_{\text{o-c}}$ is 2.3 kcal/mol47. Error bars for the unbiased and biased REST2 simulations were obtained by calculating standard deviations across blocks of 100 ns. For DDPM, error bars correspond to the standard deviation calculated across four sets of 1000000 generated points, obtained from the model trained on the same dataset.
@@
We then trained a DDPM following the procedure explained in the previous section, learning the joint probability distribution along the four collective variables that were shown to be relevant for the open/close process ($\psi_{181}$, $\phi_{182}$, $d_{181-112}$, $d_{181-221}$) and the rescaled potential energy. Although DDPM is able to generate new samples in the transition state region at the reference temperature (Figure  6B), it is not enough to resolve the transition and obtain a converged estimate of the free-energy barrier separating the two basins. Moreover, DDPM does not solve the issue of the relative stabilization of the closed conformation (Table 2). Interestingly, for the highest temperature replica, where the free-energy barrier is expected to be lower (Figure  6C), DDPM is able to resolve the barrier (Figure  6D).
Given the inability of the reference-temperature, unperturbed replica to effectively sample the transition state region between the two basins—while the most perturbed replica more uniformly explores the $\psi_{181}$ and $\phi_{182}$ dihedral angles, previously identified as the slow collective variables in loop closing/opening—we leveraged this probability distribution to bias exploration across all replicas in a new set of REST2 simulations.
Figure 7: Biasing REST2 simulations and refining with DDPM. (A) Two-dimensional bias applied to the reference temperature replica, computed following equation 6. (B) Free-energy surface obtained from the biased REST2 simulation, analyzing the reference temperature replica. (C) Prediction from DDPM trained with the biased REST2 simulations. DDPM is able to predict a second transition path, unseen in both the biased REST2 simulations and the KDE.
@@
where $V_{i}\left(\psi_{181},\phi_{182}\right)$ is the bias applied along the torsional angles $\psi_{181}$ and $\phi_{182}$, $R$ is the gas constant, $T$ is the reference temperature (300 K), and $P_{high}^{KDE}(\psi_{181},\phi_{182})$ is the joint probability distribution of these angles, estimated using kernel density estimation (KDE) 50, 51 for the high-temperature replica. The bias for each replica (see Figure  7A for the bias applied to the reference replica) is thus the negative of the free-energy profile computed along $\psi_{181}$ and $\phi_{182}$ at the highest temperature, rescaled by $\frac{\lambda_{i}}{\lambda_{high}}$ to account for the higher potential energy barriers at lower temperatures due to the REST2 rescaling scheme.
@@
Short REST2 simulations were then performed using a similar setup as the one described before but incorporating these biases, in the spirit of importance sampling. The presence of these biases significantly improved the exploration of the FES along $\psi_{181}$ and $\phi_{182}$. However, the reference replica still failed to fully sample the transition regions between the open and closed basins (Figure 7B). By retraining and applying a diffusion probabilistic model (DDPM) based on these data, we obtained an FES at the reference temperature that revealed a clear transition pathway with an estimated barrier of approximately 11-12 kcal/mol (Figure 7C). Notably, a KDE-based FES constructed from the reference replica alone yielded inaccurate transition barrier estimates, significantly exceeding those obtained from DDPM (14-15 kcal/mol, Figure S3). While methodological differences (as well as minor differences between the employed forcefields) between this work and the previous reference study31 could explain minor discrepancies—especially since the previous path was computed in a four-dimensional space incorporating $\psi_{181}$, $\phi_{182}$, $d_{181-112}$, and $d_{181-221}$—it is remarkable that both approaches converged to a nearly identical barrier of 11-12 kcal/mol31.
@@
We note that in addition to the barrier, the biased REST2 simulation now provide better estimates of the free-energy difference between the open and closed conformations, in agreement with previous computational studies (Table 2).
@@
We conducted a committor analysis along this trajectory (Figure S4) to confirm that the found reaction pathway actually represents a true transition path. Additionally, while the transition region is not perfectly resolved, an alternative pathway with a similar barrier in the direction of decreasing $\psi_{181}$ values is visible in Figure 7C. In contrast to the other pathway, committor analysis revealed that a fraction of trajectories initiated along this pathway failed to reach either the open or closed states, instead becoming trapped in an intermediate region of the FES. This suggests that this alternative pathway may involve additional slow-relaxing variables, incurring extra free-energy and/or kinetic costs.
@@
The combined use of REST2, DDPM, and importance sampling provides robust estimates of the FES along these slow modes. To further validate our approach, we repeated the simulations using the same biases but starting from the open enzyme conformation. Applying the same protocol, we identified a similar transition pathway, demonstrating that our method reliably captures the thermodynamic and mechanistic characteristics of the open/close motion (Figure 8A). In particular, the 1D projection of the projection path reveals very identical barriers (Figure 8B), with minor differences < 1 kcal/mol regarding the free-energy differnence between the two conformations (Figure 8B and Table 2).

### v2

@@ 2.9 Comparison with TREM and US
We now provide a critical comparison between the results of DDPM-REST2 and that of DDPM-TREM using a similar simulation setup. We performed TREM simulations using 32 replicas of 600 ns ranging from 279 K to 600 K, yielding an exchange rate of 0.13 (this illustrates the much lower performance of TREM compared to REST2 despite three times more replicas).
Figure 3: Comparison of two-dimensional free-energy surfaces computed along RMSD and $R_{g}$, obtained from TREM (A), TREM refined by DDPM (B), REST2 (C), and REST2 refined by DDPM (D), as well as the free-energy values along the minimum free-energy path as predicted by DDPM (E). White lines on (B) and (D) show the minimum free-energy path. (F) One-dimensional free energy profile computed along the RMSD direction, comparing the training data from TREM and REST2 (dashed lines, orange and blue, respectively) with the DDPM predictions (solid lines, orange and blue, respectively), and Umbrella Sampling as a benchmark (solid green line). Representative structures of basins $\alpha$, $\beta$, $\gamma$, and $\delta$ are shown in Figure S1.
By projecting the probability distributions estimated from the unscaled replica along (Figure 3A and C), it is clear that both TREM and REST2 achieve comparable sampling performances. Key free-energy minima are identified, corresponding to radically different conformations (a few examples numbered $\alpha$ to $\gamma$ are shown in Figure S1). We observe that while the basins corresponding to regions $\alpha$ and $\beta$ are well sampled, there are significantly fewer samples in $\gamma$ and $\delta$, and the transition regions are not well refined. As a consequence, once cannot clearly estimate the minimum free-energy paths and barriers connecting the apparent basins, in particular $\alpha$, $\beta$ and $\gamma$ on this two-dimensional surface.
When post-processing the results of the simulations with DDPM, and then use it to generate new samples for the unscaled replica conditions, much more converged free-energy surfaces can be recovered (Figure 3B and D). In particular, the transition states regions connection $\alpha$, $\beta$ and $\gamma$ are now fully resolved. These are extrapolated from the model based on higher temperature/lower potential simulations learned from the model and extrapolated in the unperturbed replica. Indeed, the corresponding areas appear sampled when climbing the replica ladder (Figure S2).
@@
Free-energy difference TREM DDPM-TREM REST2 DDPM-REST2
$F_{\beta}-F_{\alpha}$(kcal/mol) 1.11 $\pm$ 0.34 1.05 $\pm$ 0.33 1.23 $\pm$ 0.24 1.10 $\pm$ 0.24
$F_{\gamma}-F_{\beta}$(kcal/mol) 0.81 $\pm$ 0.36 0.72 $\pm$ 0.12 0.64 $\pm$ 0.30 0.81 $\pm$ 0.50
Table 1: Total free energy differences between basins $\alpha$ , $\beta$ and $\gamma$ computed from TREM, DDPM-TREM, REST2 and DDPM-REST2. Error bars for the TREM and REST2 simulations were obtained by calculating standard deviations across blocks of 100 ns, and those for DDPM were obtained by training the model on these different blocks.
Although there is good agreement between the two replica exchange strategies combined with DDPM, there is no guarantee that they actually converge toward the actual ground-truth reference of the associated conformational change. Having determined from 2D-projections that the key bottleneck to conformational exploration was (in the chosen CV-space) transitions along the RMSD to the crystal structure, we also performed US simulations along this coordinate to estimate a ground-truth 1D FES to which we could compare the results of replica exchange simulations augmented with DDPM. We stress that these data are completely unrelated to the (replica-exchange) data used to train the DDPM model.
For RMSD values below 3 Å, there is good agreement between the DDPM predictions trained on both TREM and REST2, and the umbrella sampling simulations. We note that in the one-dimensional projection, the incomplete and sparse description of barriers, visible in Figure 3A and C, is much less problematic, and both REST2 and TREM readily provide reasonable estimates.
For RMSD larger than 3 Å, the agreement between raw data from replica-exchange, DDPM, and US, is less accurate. This is due to the fact that high RMSD values correspond to a large number of possible unfolded, disordered conformations of CLN025 that are difficult to sample in all approaches (US simulations are also unlikely to be converged in that region). In particular, this indicates that the sampling provided by the replicas is not fully converged in this region of phase space. Indeed, sampling is pretty sparse and noisy in these high-RMSD regions, and even so for the highest temperature replica (Figure S2).
2.10 Comparison with long, unbiased MD trajectories
[…]
@@ 2.11 Importance of exchanges for reliable reference data
The fact that DDPM results remain sensitive to the quality of sampling in the replica exchange simulations is further illustrated by studying the impact of exchanges among replicas on the obtained results. Strikingly, it has been argued in the original DDPM-TREM formulation24, as well as in subsequent work28, that DDPM could perform well even if exchanges were vanishingly small (smaller than a few %, or even non existent).
@@
As seen in these plots, exchanges are crucial for converging the melting curve, which is a measure of the convergence of the thermodynamic quantities, and they significantly impact DDPM predictions. For instance, DDPM trained on TREM without exchanges, starting with unfolded structures, cannot generate folded structures because no folding is observed in the TREM without exchanges. Starting from a folded structure, TREM simulations can generate unfolded structures in the high-temperature replicas, but the melting curves are very noisy and DDPM alone does not improve the sampling.
@@
The main difference with previous work is that (i) here exchanges are crucial to drive chignolin folding/unfolding, whereas some of the systems investigated previously did not exhibit folded/unfolded states24; and (ii) in other more complex cases28, configurations are mixed across the replica ladder and periodically updated, therefore mitigating the effects of small exchange rates.
This highlights the advantage of using REST2 for training DDPM, as REST2 requires much fewer replicas compared to TREM. Regarding computational cost, the training of DDPM and data generation for REST2 and TREM in these examples were similarly affordable from a computational perspective (less than a dozen gpu.hrs). The gain in computational efficiency comes from the better scaling of REST2 with system size: we needed only 12 replicas with REST2, compared to 32 with TREM, achieving better exchange rates with REST2 (0.51) than with TREM (0.13) and yielding comparable sampling at the reference temperature (Figure 3A and C). This resulted in a 2.7-fold reduction in computational cost for REST2 compared to TREM for the same sampling efficiency. The gain would be even more obvious for a larger biomolecule that would require many more replicas in TREM.
@@ 2.12 Resolving large conformational changes with high activation barriers
We thus chose to focus on the loop opening/closing motion in the phosphatase PTP1B. PTP1B is a well-studied member of the Protein Tyrosine Phosphatase (PTP) superfamily, which is involved in multiple cellular processes such as glucose uptake, and proliferation 51. It is also a drug target for several diseases, including diabetes and cancer 52.
@@
Before we focus on our results, we now provide some context about this enzyme (Figure 5A). PTP1B catalyzes the hydrolysis of phosphorylated tyrosine in a two-step process involving a thiophosphate enzyme intermediate. The first step consists of a nucleophilic attack by a cysteine residue on the phosphorylated tyrosine, breaking the phosphate bond. The second step involves the hydrolysis of the intermediate through a nucleophilic attack by a water molecule. Both steps require the assistance of an aspartic residue (Asp181) which acts as a proton donor and acceptor during the reaction. Asp181 belongs to the WPD-loop, a flexible loop comprising a dozen residues (177 to 188) that exists in two conformations: the hydrolysis-incompetent open state, and the hydrolysis-competent closed state (Figure 5B). In the open state, Asp181 forms a salt-bridge with arginine Arg112, while in the closed state, Asp181 rotates to form a salt-bridge with Arg221, allowing Asp181 to get closer to the catalytic site for proton transfers.
@@
Starting first from a closed conformation (we provide further below the comparison with simulations starting from the open conformation), we propagated the simulations for 600 ns and analyzed the converged part of the trajectories. We computed the free-energy surface along the two torsional angles $\psi_{181}$ and $\phi_{182}$ whose changes allow the rotation of the Asp181-Phe182 peptide group, which have been shown to be responsible for the high free-energy barrier characterizing the transition 36. The FES (Figure 6A) is characterized by two basins, corresponding to the closed (upper left) and open (bottom right) conformations; no transition path is sampled to go from the closed to the open conformation. In addition, the free-energy difference between the two basins is in favor of the closed conformation, at odds with previous investigations and experimental results (Table 2).
Method CV-bias Starting conformation $F_{\text{open}}-F_{\text{closed}}$ (kcal/mol)
REST2 no closed 0.70 $\pm$ 0.24
DDPM-REST2 no closed 0.87 $\pm$ 0.25
REST2 yes closed 0.80-0.80 $\pm$ 1.08
DDPM-REST2 yes closed 0.92-0.92 $\pm$ 1.05
DDPM-REST2 yes open 1.75-1.75 $\pm$ 1.26
Table 2: Free energy differences between closed and open conformations, computed from REST2 starting closed, DDPM-REST2, REST2+CV-bias starting closed, REST2+CV-bias+DDPM starting closed, and REST2+DDPM+CV-biasias starting open. The experimental value for $F_{\text{open}}-F_{\text{closed}}$ is 2.3-2.3 kcal/mol53. Error bars for the unbiased and biased REST2 simulations were obtained by calculating standard deviations across blocks of 100 ns and those for DDPM were obtained by training the model on these different blocks.
@@
We then trained a DDPM following the procedure explained in the previous section, learning the joint probability distribution along the four collective variables that were shown to be relevant for the open/close process ($\psi_{181}$, $\phi_{182}$, $d_{181-112}$, $d_{181-221}$) and the rescaled potential energy. DDPM is able to generate new samples in the transition state region at the reference temperature (Figure  6B), to better resolve the open conformation basin (bottom right corner) and to explore a new conformational basin located around (360, 50). However, it is not enough to resolve the transition and obtain a converged estimate of the free-energy barrier separating the two basins. Moreover, DDPM does not solve the issue of the relative stabilization of the closed conformation (Table 2). Interestingly, for the highest temperature replica, where the free-energy barrier is expected to be lower (Figure  6C), DDPM is able to resolve the barrier (Figure  6D), which is not the case from the raw data from this replica (Figure  6C).
Given the inability of the reference-temperature, unperturbed replica to effectively sample the transition state region between the two basins (while the most perturbed replica more uniformly explores the $\psi_{181}$ and $\phi_{182}$ dihedral angles, previously identified as the slow collective variables in loop closing/opening), we leveraged this probability distribution to bias exploration across all replicas in a new set of REST2 simulations.
Figure 7: Prediction from DDPM trained on biased REST2 simulations initiated from the closed conformation. (A) Two-dimensional bias applied to the reference temperature replica, computed following equation 6. (B) Free-energy surface obtained from the biased REST2 simulation, analyzing the reference temperature replica. (C) Prediction from DDPM trained with the biased REST2 simulations. DDPM is able to predict a second transition path, unseen in both the biased REST2 simulations and the KDE.
@@
where $V_{i}\left(\psi_{181},\phi_{182}\right)$ is the bias applied along the torsional angles $\psi_{181}$ and $\phi_{182}$, $R$ is the gas constant, $T$ is the reference temperature (300 K), and $P_{high}^{KDE}(\psi_{181},\phi_{182})$ is the joint probability distribution of these angles, estimated using kernel density estimation (KDE) 56, 57 for the high-temperature replica. The bias for each replica (see Figure  7A for the bias applied to the reference replica) is thus the negative of the free-energy profile computed along $\psi_{181}$ and $\phi_{182}$ at the highest temperature, rescaled by $\frac{\lambda_{i}}{\lambda_{high}}$ to account for the higher potential energy barriers at lower temperatures due to the REST2 rescaling scheme.
@@
Short REST2 simulations were then performed using a similar setup as the one described before but incorporating these biases, in the spirit of importance sampling. The presence of these biases significantly improved the exploration of the FES along $\psi_{181}$ and $\phi_{182}$. However, the reference replica still failed to fully sample the transition regions between the open and closed basins (Figure 7B). By retraining and applying DDPM on these data, we obtained a FES at the reference temperature that reveals a clear transition pathway with an estimated barrier of approximately 11-12 kcal/mol (Figure 7C, pathway in red). Notably, a KDE-based FES constructed from the reference replica alone yielded inaccurate transition barrier estimates, significantly exceeding those obtained from DDPM (14-15 kcal/mol, Figure S4). While methodological differences (as well as minor differences between the employed forcefields) between this work and the previous reference study36 could explain minor discrepancies, especially since the previous path was computed in a four-dimensional space incorporating $\psi_{181}$, $\phi_{182}$, $d_{181-112}$, and $d_{181-221}$, it is remarkable that both approaches converged to a nearly identical barrier of 11-12 kcal/mol36.
@@
We note that in addition to the barrier, the biased REST2 simulation now provide better estimates of the free-energy difference between the open and closed conformations, in agreement with previous (unrelated) computational studies (Table 2). We stress again that this previous data has obviously not been used to train the DDPM model.
We conducted a committor analysis along this trajectory (Figure S5) to confirm that the found reaction pathway actually represents a true transition path (Figure S5). In particular, trajectories initiated around the corresponding transition state (initial configurations were selected along the REST2 trajectories) fall in either the closed or open basin, with a ratio close to 50:50 at the transition state location (Figure S5).
@@
Additionally, while the transition region is not perfectly resolved, an alternative pathway with a similar barrier in the direction of decreasing $\psi_{181}$ values is visible in Figure 7C, passing through a transition station located around (120, -50). While it is not perfectly resolved starting from the closed state, it’s quite unambiguous when starting from the open conformation (see below). In contrast to the other pathway, a committor analysis reveals that a fraction of trajectories initiated along this pathway fail to reach either the open or closed states, instead becoming trapped in an intermediate region of the FES. This suggests that this alternative pathway may involve additional slow-relaxing variables, incurring extra free-energy and/or kinetic costs.
@@
The combined use of REST2, DDPM, and importance sampling provides robust estimates of the FES along what are indeed slow modes of the loop motion. We finally verified that the replica-exchange simulations were converged, in that sense that they were not sensitive to the starting structure. We thus repeated the simulations using the same biases but starting from the open loop conformation. Applying the same protocol, we identified a similar transition pathway, demonstrating that our method reliably captures the thermodynamic and mechanistic characteristics of the open/close motion (Figure 8A). In particular, the 1D projection of the projection path reveals very identical barriers (Figure 8B), with minor differences < 1 kcal/mol regarding the free-energy difference between the two conformations (Figure 8B and Table 2).